# Adsorption of volatiles on dust grains in protoplanetary disks

### v2

@@ 3.1 Setups of KMC simulations
3.2 Surface coverage conditions simulated by KMC
@@
Simulations are conducted for dust grains situated in the mid-plane of a model PPD, employing density and temperature profiles consistent with a young stellar object of bolometric luminosity $L\sim 4L_{\odot}$. The temperature profile follows $T\propto R^{-0.4}$ due to surface flaring effects 1997ApJ...490..368C, with specific forms given by:
@@
Across the radial domain of $0.1<(R/\mathrm{AU})<200$, KMC simulations revealed distinct surface coverage patterns for carbonaceous surfaces and the $(001)^{-}$ silicate surface, as illustrated in Figures 5. Silicate surfaces undergo a transition from bare to coated states at $R\lesssim 0.2~\mathrm{AU}$, a consequence of the substantial chemisorption energies characteristic of $\mathrm{CO}$ and $\mathrm{H_{2}O}$ on these substrates. In contrast, KMC results clearly identify a snowline on grain surfaces for water on carbonaceous grains at $R\simeq 8~\mathrm{AU}$ ($T\simeq 120~{\rm K}$), interior to which surfaces remain uncoated due to elevated thermal energies (Figure 5). This snowline represents a distinct concept from the conventional snowline marking $\mathrm{H_{2}O}$ ice condensation at $T\sim 170~{\rm K}$ (2000ApJ...528..995S), with the former being more directly relevant to substrate material coagulation processes. Employing the previously assumed adsorption energy of $\simeq 0.45~{\rm eV}$ would reposition the $\mathrm{H_{2}O}$ snowline at $R\simeq 2~\mathrm{AU}$.
@@
The fundamental differences in adsorption behavior between carbonaceous and silicate substrates have likely remained obscured in previous research due to systematic misinterpretations of temperature-programmed desorption (TPD) experimental data. This misinterpretation stems from the complex nature of water cluster formation on substrate surfaces. Within water molecule clusters, each molecule typically participates in two hydrogen bonds on average, mirroring the structural configuration of hexagonal ice crystals (Ih). This hydrogen bonding network contributes an internal binding energy of approximately $\epsilon_{\mathrm{hb}}\sim 0.42~\mathrm{eV}$ per molecule (where the subscript "hb" denotes hydrogen bonds; see also silverstein2000strength).
@@
Conventional TPD experiments are conducted under conditions substantially different from astrophysical environments: temperatures well below the triple point of water and gas-phase molecule number densities orders of magnitude higher than those found in protoplanetary disks or the interstellar medium. These experimental conditions promote the formation of extensive water clusters on substrate surfaces PhysRevB.79.235440; PhysRevB.84.033402; bjorneholm2016water. Consequently, molecular addition or removal processes during TPD measurements primarily occur at the interfaces of these water clusters rather than at the actual substrate-adsorbate boundary. This clustering phenomenon could lead to a possible misinterpretation of TPD data: the slopes observed in Arrhenius plots actually reflect $\epsilon_{\mathrm{hb}}/k_{\mathrm{B}}$, the energy scale associated with hydrogen bond breaking, rather than the true substrate adsorption energy $\Delta\epsilon_{\mathrm{ad}}/k_{\mathrm{B}}$. For silicate substrates, the situation is further complicated by the exceptional strength of chemisorption bonds, which remain intact at typical protoplanetary disk temperatures ($T\lesssim 10^{3}~{\rm K}$). Under these conditions, any molecular desorption processes are governed by hydrogen bond rupture within water clusters, again causing TPD measurements to yield $\epsilon_{\mathrm{hb}}$ rather than the fundamental substrate adsorption energy. Accurate interpretation of adsorption mechanisms and binding energies is therefore essential for reliable modeling of astrophysical processes involving dust grains.
4 Discussions
@@
We have combined vdW-corrected ab-initio DFT calculations and kinetic Monte Carlo simulations to quantify how key volatiles ($\mathrm{H_{2}O}$, $\mathrm{CO}$, and $\mathrm{H_{2}}$) adsorb onto carbonaceous and silicate grain surfaces, and how these material-dependent adsorption energies, combined with KMC simulations, translate into different surface coverage patterns. The results reveal a clear dichotomy between weak physisorption on carbonaceous grains and strong chemisorption on silicates, with important consequences for grain growth and volatile partitioning.
@@
4.1 Dependence of snowline locations on
@@
Consequently, history-dependent snowline locations could vary in different disks, occurring at radii corresponding to different temperatures. Simply inferring their locations based on a fixed temperature (e.g., $20~{\rm K}$ for CO) may therefore lead to mismatches. Dust rings and gaps have been found to be nearly ubiquitous in disks. The prevalent hypothesis for their origin is that they are caused by planet–disk interactions (dong15gap; dong17doublegap; dong17gap), while the snowline hypothesis (zhang15) has been questioned due to mismatches between the observed ring/gap locations and the expected snowline locations (long18; bae18; vandermarel19). Re-examining the connection between the two, while accounting for each disk’s thermal and dynamical history, is therefore needed.
4.2 Impact of multi-species adsorption on CO surface snowlines and gas depletion
The analysis of surface coverage conditions reveals that, based on the disk temperature and hydrodynamic profiles, $R\sim 20~\mathrm{AU}$ approximately divides the grain surface snowline for $\mathrm{CO}$ on amorphous carbon grains, corresponding to a local temperature of $T\simeq 80~{\rm K}$ according to the disk temperature profile defined in equation (3). This temperature appears anomalously high when compared to existing experimental and computational studies, which typically report CO condensation temperatures around $\sim 20~{\rm K}$ (2017ApJ...849...80C). However, detailed examination of Figure 5 indicates that CO condensation initiation is not an isolated process but rather occurs through coordinated condensation with $\mathrm{H_{2}O}$ molecules.
@@ 4.2 Impact of multi-species adsorption on CO surface snowlines and gas depletion
4.3 Molecule Condensation and Gas-Phase Depletion
Our calculations indicate that the CO midplane snowline lies significantly closer to the central star than conventional estimates suggest. In a representative PPD around a solar-type star with $T(1~\mathrm{AU})\simeq 150~{\rm K}$, the CO snowline is expected at around $30~\mathrm{AU}$ (e.g., Oberg2023). When adopting the higher midplane temperature used in our disk model ($280~{\rm K}$ instead of $150~{\rm K}$ at $1~\mathrm{AU}$), this discrepancy becomes even more pronounced. This inward shift of CO snowline in our model mainly arises from enhanced $\mathrm{CO}$$\mathrm{H_{2}O}$ binding, which promotes earlier condensation of CO onto water-rich grain surfaces.
Observationally, molecular snowline locations can be inferred from chemical tracers of gas-phase abundances, such as $\mathrm{N_{2}H^{+}}$ for CO, which usually place the CO snowline beyond 40 AU in disks around solar-type stars (Qi2013; Qi2019). However, these diagnostics are not always uniquely linked to volatile depletion at the corresponding snowline and can be affected by optical depth effects, non-thermal desorption, or chemical processing (van'tHoff2017A&A...599A.101V). These measurements typically trace the traditional $20~{\rm K}$ condensation front, where the CO gas abundance drops dramatically. In contrast, our models predict the onset of CO condensation significantly closer to the star, at a region not directly captured by such observations. This suggests that a substantial disk region may contain partially CO-ice-coated grains, effectively extending the CO-ice zone inward compared to classical estimates.
@@ 4.3 Molecule Condensation and Gas-Phase Depletion
Our models also predict that $\mathrm{H_{2}O}$ ice can sublimate at lower temperatures and thus at larger disk radii, when it is coated on carbonaceous grains. These shifts of snowline locations have important implications for interpreting the inner disk chemistry revealed by JWST MIR spectra. The initial gas-phase C/O ratio in disks is largely determined by the relative positions of major volatile snowlines. For example, it is expected to be lower inside the $\mathrm{H_{2}O}$ snowline than in the region between $\mathrm{H_{2}O}$ and $\mathrm{CO}$ snowlines (Oberg11). Current interpretations of JWST spectra often invoke the sublimation of icy grains as they drift inward and cross various snowlines (e.g., Mah2023). Recent models that include CO entrapment in water ice suggest elevated C/O and C/H ratio inside the water snowline, by up to a factor of 10 over $1~{\rm Myr}$ (Williams2025). This process is analogous to the closer-in onset of CO condensation predicted in our calculations. Precise knowledge of snowline locations and their temperature dependencies is thus essential for modeling disk chemistry. The outward shift of the water snowline predicted here adds an additional layer of complexity that future chemo-dynamical models should consider. Moreover, the lower sublimation temperature of water ice implies that $\mathrm{H_{2}O}$ vapor may persist at colder temperatures, motivating future far-infrared observations with facilities such as PRIMA to probe water beyond the canonical snowline.
5 Summary
@@
We have combined vdW-corrected ab-initio DFT calculations (§2) and kinetic Monte Carlo simulations (§3) to quantify how key volatiles ($\mathrm{H_{2}O}$, $\mathrm{CO}$, and $\mathrm{H_{2}}$) adsorb onto carbonaceous and silicate grain surfaces. The resulting, material-dependent adsorption energies, when propagated through the KMC framework, produce distinct surface coverage patterns and well-defined snowlines on grain surfaces in a representative protoplanetary disk model (Figure 5). Our calculations also demonstrate a clear dichotomy between weak physisorption on carbonaceous grains and strong chemisorption on silicates (Tables 1 and 2), which causes carbonaceous grains to lose their volatile coatings in the inner disk while silicates remain molecularly coated over a much wider range of radii.
@@
Building on this physical picture, we show that history-dependent initial surface conditions and evolutionary paths can shift snowline locations (§4.1), that multi-species adsorption and enhanced $\mathrm{CO}$ binding in $\mathrm{H_{2}O}$-rich ices can move the effective CO surface snowline inward and strengthen gas-phase CO depletion (§4.2), and that these effects together impact disk gas-mass estimates, the interpretation of snowline-sensitive tracers, and JWST spectra in planet-forming disks (§4.3).
@@ Appendix A Thermodynamic Corrections for Adsorption Energies
The thermodynamic treatment of adsorbate-substrate systems requires careful consideration of their inherent complexity compared to isolated molecular systems. Dust grain substrates possess substantially larger heat capacities than individual adsorbate molecules and maintain stiff thermal equilibrium with the diffuse radiation fields characteristic of interstellar media and protoplanetary disks 1997ApJ...490..368C. Consequently, these substrates effectively function as isothermal heat reservoirs during adsorption-desorption processes. When molecular desorption occurs, the thermal relaxation timescale in the gas phase through collisions with other molecules can be estimated following DraineBook:
@@ B.2 Amorphous Carbon Substrate Construction
The final optimized structure possessed a supercell with transverse dimensions of $(9.09~\text{\AA })^{2}$ and exhibited a mixed hybridization state distribution: 47% of carbon atoms in $sp^{2}$ configuration and 53% in $sp^{3}$ configuration. The resulting carbon bulk demonstrated an average thickness of $\sim 5.1$$\text{\AA }$ and a density of $\sim 2.4~{\rm g}~{\rm cm}^{-3}$, values consistent with experimentally characterized amorphous carbon materials. Surface hydrogen coverage reached $\sim 0.16~{\rm cm}^{-2}$, corresponding to approximately 1% atomic hydrogen concentration for a spherical grain of radius $0.05~{\rm\mu m}$, in excellent agreement with experimental measurements 1998A&A...332..291J.
@@ Appendix C General procedures of KMC simulations
KMC simulations have been performed extensively for various studies for surface physical and chemical properties (e.g. 2019FrCh....7..202A). Three distinct processes are modeled to occur on the simulated grain surfaces, each characterized by specific rate equations:

### v3

@@ III.1 Setups of KMC simulations
III.2 Surface coverage conditions simulated by KMC
@@
Simulations are conducted for dust grains situated in the mid-plane of a model PPD, employing density and temperature profiles consistent with a young stellar object of bolometric luminosity $L\sim 4L_{\odot}$. The temperature profile follows $T\propto R^{-0.4}$ due to surface flaring effects Chiang and Goldreich (1997), with specific forms given by:
@@
Across the radial domain of $0.1<(R/\mathrm{AU})<200$, KMC simulations revealed distinct surface coverage patterns for carbonaceous surfaces and the $(001)^{-}$ silicate surface, as illustrated in Figures 5. Silicate surfaces undergo a transition from bare to coated states at $R\lesssim 0.2~\mathrm{AU}$, a consequence of the substantial chemisorption energies characteristic of $\mathrm{CO}$ and $\mathrm{H_{2}O}$ on these substrates. In contrast, KMC results clearly identify a snowline on grain surfaces for water on carbonaceous grains at $R\simeq 8~\mathrm{AU}$ ($T\simeq 120~{\rm K}$), interior to which surfaces remain uncoated due to elevated thermal energies (Figure 5). This snowline represents a distinct concept from the conventional snowline marking $\mathrm{H_{2}O}$ ice condensation at $T\sim 170~{\rm K}$ (Sasselov and Lecar, 2000), with the former being more directly relevant to substrate material coagulation processes. Employing the previously assumed adsorption energy of $\simeq 0.45~{\rm eV}$ would reposition the $\mathrm{H_{2}O}$ snowline at $R\simeq 2~\mathrm{AU}$.
@@
The fundamental differences in adsorption behavior between carbonaceous and silicate substrates have likely remained obscured in previous research due to systematic misinterpretations of temperature-programmed desorption (TPD) experimental data. This misinterpretation stems from the complex nature of water cluster formation on substrate surfaces. Within water molecule clusters, each molecule typically participates in two hydrogen bonds on average, mirroring the structural configuration of hexagonal ice crystals (Ih). This hydrogen bonding network contributes an internal binding energy of approximately $\epsilon_{\mathrm{hb}}\sim 0.42~\mathrm{eV}$ per molecule (where the subscript “hb” denotes hydrogen bonds; see also Silverstein et al. (2000)).
@@
Conventional TPD experiments are conducted under conditions substantially different from astrophysical environments: temperatures well below the triple point of water and gas-phase molecule number densities orders of magnitude higher than those found in protoplanetary disks or the interstellar medium. These experimental conditions promote the formation of extensive water clusters on substrate surfaces Leenaerts et al. (2009); Ma et al. (2011); Björneholm et al. (2016). Consequently, molecular addition or removal processes during TPD measurements primarily occur at the interfaces of these water clusters rather than at the actual substrate-adsorbate boundary. This clustering phenomenon could lead to a possible misinterpretation of TPD data: the slopes observed in Arrhenius plots actually reflect $\epsilon_{\mathrm{hb}}/k_{\mathrm{B}}$, the energy scale associated with hydrogen bond breaking, rather than the true substrate adsorption energy $\Delta\epsilon_{\mathrm{ad}}/k_{\mathrm{B}}$. For silicate substrates, the situation is further complicated by the exceptional strength of chemisorption bonds, which remain intact at typical protoplanetary disk temperatures ($T\lesssim 10^{3}~{\rm K}$). Under these conditions, any molecular desorption processes are governed by hydrogen bond rupture within water clusters, again causing TPD measurements to yield $\epsilon_{\mathrm{hb}}$ rather than the fundamental substrate adsorption energy. Accurate interpretation of adsorption mechanisms and binding energies is therefore essential for reliable modeling of astrophysical processes involving dust grains.
IV Discussions
@@
We have combined vdW-corrected ab-initio DFT calculations and kinetic Monte Carlo simulations to quantify how key volatiles ($\mathrm{H_{2}O}$, $\mathrm{CO}$, and $\mathrm{H_{2}}$) adsorb onto carbonaceous and silicate grain surfaces, and how these material-dependent adsorption energies, combined with KMC simulations, translate into different surface coverage patterns. The results reveal a clear dichotomy between weak physisorption on carbonaceous grains and strong chemisorption on silicates, with important consequences for grain growth and volatile partitioning. We notice that, although crystalline enstatite is employed as a proxy for more complex amorphous silicates, the fact that chemisorption energies on silicate surfaces are an order of magnitude stronger than on carbonaceous surfaces remains relatively robust: the adsorption energy magnitudes ($|\Delta\epsilon_{\rm ad}|\gtrsim 1~{\rm eV}$ for $\mathrm{H_{2}O}$ and $\sim 0.5-1~{\rm eV}$ for CO) are not significantly influenced by the crystalline nature of the substrate (Tables 1 and 2).
@@
IV.1 Dependence of snowline locations on
@@
Consequently, history-dependent snowline locations could vary in different disks, occurring at radii corresponding to different temperatures. Simply inferring their locations based on a fixed temperature (e.g., $20~{\rm K}$ for CO) may therefore lead to mismatches. Dust rings and gaps have been found to be nearly ubiquitous in disks. The prevalent hypothesis for their origin is that they are caused by planet–disk interactions (Dong et al., 2015, 2017; Dong and Fung, 2017), while the snowline hypothesis (Zhang et al., 2015) has been questioned due to mismatches between the observed ring/gap locations and the expected snowline locations (Long et al., 2018; Bae et al., 2018; van der Marel et al., 2019). Re-examining the connection between the two, while accounting for each disk’s thermal and dynamical history, is therefore needed.
IV.2 Impact of multi-species adsorption on CO surface snowlines and gas depletion
The analysis of surface coverage conditions reveals that, based on the disk temperature and hydrodynamic profiles, $R\sim 20~\mathrm{AU}$ approximately divides the grain surface snowline for $\mathrm{CO}$ on amorphous carbon grains, corresponding to a local temperature of $T\simeq 80~{\rm K}$ according to the disk temperature profile defined in equation (3). This temperature appears anomalously high when compared to existing experimental and computational studies, which typically report CO condensation temperatures around $\sim 20~{\rm K}$ (Cazaux et al., 2017). However, detailed examination of Figure 5 indicates that CO condensation initiation is not an isolated process but rather occurs through coordinated condensation with $\mathrm{H_{2}O}$ molecules.
@@ IV.2 Impact of multi-species adsorption on CO surface snowlines and gas depletion
IV.3 Molecule Condensation and Gas-Phase Depletion
Our calculations indicate that the CO midplane snowline lies significantly closer to the central star than conventional estimates suggest. In a representative PPD around a solar-type star with $T(1~\mathrm{AU})\simeq 150~{\rm K}$, the CO snowline is expected at around $30~\mathrm{AU}$ (e.g., Öberg et al., 2023). When adopting the higher midplane temperature used in our disk model ($280~{\rm K}$ instead of $150~{\rm K}$ at $1~\mathrm{AU}$), this discrepancy becomes even more pronounced. This inward shift of CO snowline in our model mainly arises from enhanced $\mathrm{CO}$$\mathrm{H_{2}O}$ binding, which promotes earlier condensation of CO onto water-rich grain surfaces.
Observationally, molecular snowline locations can be inferred from chemical tracers of gas-phase abundances, such as $\mathrm{N_{2}H^{+}}$ for CO, which usually place the CO snowline beyond 40 AU in disks around solar-type stars (Qi et al., 2013, 2019). However, these diagnostics are not always uniquely linked to volatile depletion at the corresponding snowline and can be affected by optical depth effects, non-thermal desorption, or chemical processing (van ’t Hoff et al., 2017). These measurements typically trace the traditional $20~{\rm K}$ condensation front, where the CO gas abundance drops dramatically. In contrast, our models predict the onset of CO condensation significantly closer to the star, at a region not directly captured by such observations. This suggests that a substantial disk region may contain partially CO-ice-coated grains, effectively extending the CO-ice zone inward compared to classical estimates.
@@ IV.3 Molecule Condensation and Gas-Phase Depletion
Our models also predict that $\mathrm{H_{2}O}$ ice can sublimate at lower temperatures and thus at larger disk radii, when it is coated on carbonaceous grains. These shifts of snowline locations have important implications for interpreting the inner disk chemistry revealed by JWST MIR spectra. The initial gas-phase C/O ratio in disks is largely determined by the relative positions of major volatile snowlines. For example, it is expected to be lower inside the $\mathrm{H_{2}O}$ snowline than in the region between $\mathrm{H_{2}O}$ and $\mathrm{CO}$ snowlines (Öberg et al., 2011). Current interpretations of JWST spectra often invoke the sublimation of icy grains as they drift inward and cross various snowlines (e.g., Mah et al., 2023). Recent models that include CO entrapment in water ice suggest elevated C/O and C/H ratio inside the water snowline, by up to a factor of 10 over $1~{\rm Myr}$ (Williams et al., 2025). This process is analogous to the closer-in onset of CO condensation predicted in our calculations. Precise knowledge of snowline locations and their temperature dependencies is thus essential for modeling disk chemistry. The outward shift of the water snowline predicted here adds an additional layer of complexity that future chemo-dynamical models should consider. Moreover, the lower sublimation temperature of water ice implies that $\mathrm{H_{2}O}$ vapor may persist at colder temperatures, motivating future far-infrared observations with facilities such as PRIMA to probe water beyond the canonical snowline.
V Summary
We have combined vdW-corrected ab-initio DFT calculations (§II) and kinetic Monte Carlo simulations (§III) to quantify how key volatiles ($\mathrm{H_{2}O}$, $\mathrm{CO}$, and $\mathrm{H_{2}}$) adsorb onto carbonaceous and silicate grain surfaces. Although complete representations of amorphous carbon and amorphous silicates are computationally prohibitive, the surfaces of model amorphous carbon and crystalline enstatite have been sampled as proxies. The strong chemisorption energy ($\gtrsim 1~{\rm eV}$) of $\mathrm{H_{2}O}$ and CO molecules on silicate surfaces renders the prospective error from using such proxies relatively less important. By sampling various facets of crystalline enstatite, the bracketing approach yields results for silicate surfaces that are reasonable, semi-quantitative approximations to realistic amorphous silicates.
@@ V Summary
The material-dependent adsorption energies, when incorporated into the KMC framework, lead to distinct surface coverage patterns and well-defined snowlines on grain surfaces in a representative protoplanetary disk model (Figure 5). Our calculations have also demonstrated a dichotomy between weak physisorption on carbonaceous grains and strong chemisorption on silicate grains (Tables 1 and 2), which results in carbonaceous grains losing their volatile coatings in the inner disk while silicate grains remain molecularly coated over a much wider radial range.
@@
Building on this physical picture, we show that history-dependent initial surface conditions and evolutionary paths can shift snowline locations (§IV.1), that multi-species adsorption and enhanced $\mathrm{CO}$ binding in $\mathrm{H_{2}O}$-rich ices can move the effective CO surface snowline inward and strengthen gas-phase CO depletion (§IV.2), and that these effects together impact disk gas-mass estimates, the interpretation of snowline-sensitive tracers, and JWST spectra in planet-forming disks (§IV.3).
@@ Appendix A Thermodynamic Corrections for Adsorption Energies
The thermodynamic treatment of adsorbate-substrate systems requires careful consideration of their inherent complexity compared to isolated molecular systems. Dust grain substrates possess substantially larger heat capacities than individual adsorbate molecules and maintain stiff thermal equilibrium with the diffuse radiation fields characteristic of interstellar media and protoplanetary disks Chiang and Goldreich (1997). Consequently, these substrates effectively function as isothermal heat reservoirs during adsorption-desorption processes. When molecular desorption occurs, the thermal relaxation timescale in the gas phase through collisions with other molecules can be estimated following Draine (2011):
@@ B.2 Amorphous Carbon Substrate Construction
The final optimized structure possessed a supercell with transverse dimensions of $(9.09~\text{\AA })^{2}$ and exhibited a mixed hybridization state distribution: 47% of carbon atoms in $sp^{2}$ configuration and 53% in $sp^{3}$ configuration. The resulting carbon bulk demonstrated an average thickness of $\sim 5.1$$\text{\AA }$ and a density of $\sim 2.4~{\rm g}~{\rm cm}^{-3}$, values consistent with experimentally characterized amorphous carbon materials. Surface hydrogen coverage reached $\sim 0.16~{\rm cm}^{-2}$, corresponding to approximately 1% atomic hydrogen concentration for a spherical grain of radius $0.05~{\rm\mu m}$, in excellent agreement with experimental measurements Jager et al. (1998).
@@ Appendix C General procedures of KMC simulations
KMC simulations have been performed extensively for various studies for surface physical and chemical properties (e.g. Andersen et al., 2019). Three distinct processes are modeled to occur on the simulated grain surfaces, each characterized by specific rate equations: